OpenStudy (anonymous):

I will give you a medal and be your fan! Solve for n. V=n/g A.N=v/g B.N=g/v C.N=v+g D.N=vg

OpenStudy (anonymous):

I think it's B but im not sure

OpenStudy (anonymous):

What do you think it is @Ryan88 ?

OpenStudy (anonymous):

It sounds like the answer would be A then from how you're explaining it.

OpenStudy (anonymous):

Why are you deleting everything?

OpenStudy (anonymous):

\[v = \frac{ n }{ g }\] Multiply both sides by g. \[v*g = \frac{ n }{ g }*g\] The g's on the right side of the equal sign cancel leaving what left?

OpenStudy (anonymous):

Yea, for confusion sake, I just deleted it all so I could re type it.

OpenStudy (anonymous):

So the answer is A?

OpenStudy (anonymous):

The g's on the right cancel each other out leaving : \[v*g = n\]

OpenStudy (anonymous):

So it's D

OpenStudy (anonymous):

That's right!

OpenStudy (anonymous):

I was not paying attention to the division sign,I thought it was multiplication!!! Thank you!!!!!!!!!

OpenStudy (anonymous):

No problem, sorry if I made it a little confusing with my typo earlier

OpenStudy (mkmkasim):

vg=n. since you multipy the g

OpenStudy (anonymous):

Thanks @mkmkasim but the question has been answered